OpenStudy (elusive):

How do you solve Arctan(-sqrt3) ?

OpenStudy (muratovic):

It basically asks you at which angle is tangent equal to\[-\sqrt{3}\]So you can say\[\tan \alpha=-\sqrt{3}\]I hope that you know the angle where tangent equals minus square root of three.

OpenStudy (elusive):

tan: sin/cos

OpenStudy (elusive):

so 2pi/3 or 5pi/3

OpenStudy (elusive):

????

OpenStudy (mathmale):

"How do you solve Arctan(-sqrt3)?" First of all, how much familiarity have you with inverse trig functions? \[y=\tan ^{-1}(-\sqrt{3})\]

OpenStudy (mathmale):

can be verbalized as "the angle whose tangent (ratio) is \[\frac{ -\sqrt{3} }{ 1 }\]

OpenStudy (elusive):

yes. according to the unit circle, i got 2pi/3 or 5pi/3 ,. I don't know which one to choose

OpenStudy (mathmale):

Here we're using y to represent that angle. It follows that \[\tan y = \frac{ -\sqrt{3} }{ 1 }=\frac{ opp~side }{ adj~side}\]

OpenStudy (mathmale):

Are you able to draw this situation, namely, a triangle whose opposite side, being negative, is in Quadrant III or Quadrant IV, and whose adjacent side is +1?

OpenStudy (jonathan34):



OpenStudy (jonathan34):

wrong drawling lol :O

OpenStudy (mathmale):